OpenStudy (anonymous):

f(x)=6*x^2-x f(x+h)-f(x)/h Solve the Function. (Show your work please)

OpenStudy (turingtest):

ugh.. these things are a pain to type out but I'll try. Wait do you need the limit as h-->0 or just the "f(x+h)-f(x)/h" part?

OpenStudy (anonymous):

in my prob it says that H does not equal to zero

OpenStudy (anonymous):

U still there?

OpenStudy (turingtest):

good, no limits then\[f(x)=6x^2-x\]\[f(x+h)=6(x+h)^2-(x+h)=6x^2+12xh+h^2-x-h\]\[f(x+h)-f(x)=6x^2+12xh+h^2-x-h-(6x^2-x)\]\[=12xh+h^2-h\]\[[f(x+h)-f(x)]/h=(12xh+h^2-h)/h=12x+h-1\]

OpenStudy (turingtest):

sorry takes a while to write it out

OpenStudy (anonymous):

no you're fine... i was just checking.

OpenStudy (turingtest):

PS in the future you want to write this as [f(x+h)-f(x)]/h the brackets are important to show that everything is over h, not just the f(x) part

OpenStudy (anonymous):

my bad. not the brightest crayon in the box when it comes to math. haha

OpenStudy (turingtest):

well if you got to precalculus (that's what this is right?) then you're doing fine. Parentheses can be tricky when writing on this site

OpenStudy (anonymous):

so the answer is 12x+h-1? If it is, my hw site is saying that it's wrong.

OpenStudy (turingtest):

sorry I lost a 6 it should be 12x+6h-1

OpenStudy (turingtest):

If this was calculus that would not have mattered, that's why I ignored it

OpenStudy (turingtest):

f(x)=6x^2−x f(x+h)=6(x+h^)2−(x+h)=6x^2+12xh+6h^2−x−h f(x+h)−f(x)=6x^2+12xh+6h^2−x−h−(6x^2−x) =12xh+6h^2−h [f(x+h)−f(x)]/h=(12xh+6h^2−h)/h=12x+6h−1 is how it was supposed to be

OpenStudy (anonymous):

yeah, it was the 6. this is college algebra btw.

OpenStudy (turingtest):

Oh, well FYI they are giving you this because it is essential to the definition of the derivative, which is \[\lim_{h \rightarrow 0}[f(x+h)-f(x)]/h\]you may not have done limits yet but you can see that if h was zero you would get 12x-1 whether the "h" had a 6 in front or not.

OpenStudy (anonymous):

Yeah, we just began doing these...